# Data-driven computation of adjoint sensitivities without adjoint solvers: An application to thermoacoustics

### v1

@@ III Adjoint of the Echo State Network
In this section, we derive the adjoint of the parameter-aware ESN. When the ESN runs in closed-loop, it acts as an autonomous dynamical system. This is the configuration we use for the computation of the adjoint. We consider the time-averaged objective functional (2), which we wish to optimize. In discrete time, we reformulate it as a sum over a discrete number of time steps, denoted by $N$
@@
$\mathbf{q}^{+}(i)$ $=\frac{1}{N}\frac{\partial\tilde{\mathcal{J}}(\mathbf{r}(i))}{\partial\mathbf{r}(i)}^{\top}+\frac{\partial\mathbf{r}(i+1)}{\partial\mathbf{r}(i)}^{\top}\mathbf{q}^{+}(i+1),$ (12b)
$\mathbf{q}^{+}(N)$ $=\frac{1}{N}\frac{\partial\tilde{\mathcal{J}}(\mathbf{r}(N))}{\partial\mathbf{r}(N)}^{\top}.$ (12c)
Equation (12) is in a form similar to the continuous time equations (5), however, we have replaced the system’s Jacobian with the ESN Jacobian, thus, the adjoint variables are the adjoint of the ESN’s reservoir state. Algorithmically, we first obtain the direct solution, $\mathbf{r}(i)$, as a baseline trajectory by running the ESN in closed-loop for a time window. Then, the adjoint equations (12b) are solved backwards in time starting from the terminal condition $\mathbf{q}^{+}(N)$ (12c). This procedure requires the computation of the Jacobian, $\partial\mathbf{r}(i+1)/\mathbf{r}(i)$, and the partial derivative with respect to the parameters $\partial\mathbf{r}(i)/\partial\mathbf{p}$, associated with the ESN. These gradients are provided in Appendix B.
@@
$\frac{\partial\tilde{\mathcal{J}}(\mathbf{r}(i))}{\partial\mathbf{r}(i)}=2\mathbf{W}_{out}^{\top}\mathbf{W}_{out}\mathbf{r}(i).$ (14)
@@ IV Thermoacoustics
We consider a prototypical thermoacoustic system known as the Rijke tube, which consists of a duct with a heat source located inside. The Rijke tube is a setup used to propose fundamental methods of thermoacoustic instabilities because it captures the qualitative nonlinear dynamics and bifurcations observed in real-life applications [e.g., 45, 35, 6]. In this model, we make the following assumptions; (i) the acoustics are one-dimensional, (ii) the mean-flow has a low Mach number and the acoustic variables are modelled as perturbations on top of a mean flow with uniform density across the tube, (iii) the heat release is compact, and (iv) the boundary conditions are ideal, i.e., fully reflective. The governing partial differential equations (PDEs) are derived from conservation of momentum, energy, and mass [e.g., 5]
@@ V Thermoacoustic Echo State Network
$\frac{d\mathcal{J}}{d\mathbf{p}}$ $=\sum_{i=1}^{N}\frac{\partial\tilde{\mathcal{J}}(\mathbf{r}(i))}{\partial\mathbf{r}(i)}\mathbf{Q}(i),$ (22a)
@@
$\mathbf{q}^{+}(i)$ $=\frac{1}{N}\frac{\partial\tilde{\mathcal{J}}(\mathbf{r}(i))}{\partial\mathbf{r}(i)}^{\top}+\frac{\partial\mathbf{r}(i+1)}{\partial\mathbf{r}(i)}^{\top}\mathbf{q}^{+}(i+1)+\frac{\partial\mathbf{r}(i+1+N_{\tau})}{\partial\mathbf{r}(i)}^{\top}\mathbf{q}^{+}(i+1+N_{\tau}),$ (23b)
$\mathbf{q}^{+}(N)$ $=\frac{1}{N}\frac{\partial\tilde{\mathcal{J}}(\mathbf{r}(N))}{\partial\mathbf{r}(N)}^{\top},$ (23c)
@@
The derivation of the time-delayed adjoint is provided in Appendix A. The time-delayed Jacobian is provided in Appendix B.
@@
VI Learning thermoacoustics and parametric sensitivities
@@ VI.1 Training and validation
For training and validation, we create a dataset from time-delays, $\tau=\{0.1,0.15,0.2,0.25,0.3\}$ and heat release strengths, $\beta=\{1.0,2.0,3.0,4.0,5.0\}$, giving a total of 25 thermoacoustic regimes. These parameters generate limit cycle solutions except for $(\beta=5.0,\tau=0.15)$, which produces a quasiperiodic solution. We fix the flame location $x_{f}=0.2$, modal damping coefficients $c_{1}=0.1$ and $c_{2}=0.06$, the number of Galerkin modes $n_{g}=4$, and the number of Chebyshev discretization $n_{c}=10$ [6]. With four Galerkin modes, we observe the rich nonlinear behaviour associated with thermoacoustics. We generate the data by time-marching the ODEs (17) using the scipy.odeint tool with a time step of 0.001, which we upsample to 0.01 for the ESN. We remove a transient of 200 time units, select the next 4 time units as washout and 8 time units as training data. (We found that increasing the training length further does not improve the performance whereas it increases the computational effort.) We randomly split this dataset into 20 regimes for training and 5 regimes for validation of the hyperparameters. During validation, we evaluate the closed-loop performance of 5 random realisations of the ESN, i.e., realisations of the randomly generated input and state matrices, over 5 trajectories, or folds, of length 4 time units (approximately 2 periods) starting from random time steps. As error metric, we compute relative $\ell_{2}$-error between the prediction, $\hat{\mathbf{y}}$ and the ground truth $\mathbf{y}$
@@
We consider three networks: (i) the standard parameter-aware ESN (9) with the Galerkin amplitudes as the input and output vectors $\mathbf{y}_{in}(i)=\hat{\mathbf{y}}(i)=[\hat{\mathbf{\bm{\eta}}}(i);\;\hat{\mathbf{\bm{\mu}}}(i)]$, (ii) we augment the input and output vectors with the advection variable $\mathbf{y}_{in}(i)=\hat{\mathbf{y}}(i)=[\hat{\mathbf{\bm{\eta}}}(i);\;\hat{\mathbf{\bm{\mu}}}(i);\;\hat{\mathbf{v}}(i)]$, and (iii) thermoacoustic ESN (T-ESN) (21). The standard and the advection-augmented parameter-aware ESNs have both $\beta$ and $\tau$ as explicit parameters in the input channel, i.e., in Eq. (9) $\mathbf{p}=[\beta;\;\tau]$. The T-ESN has $\beta$ as a parameter in the input channel and $\tau$ is modelled as a time-delay in the delayed velocity term in the input channel $u_{f}(t-\tau)$ (21a). For comparison, we fix the reservoir size to 1200 and connectivity of the reservoir state variables to 20. The optimal hyperparameters are provided in Table 1. The first row of Table 1 shows the ranges from which the hyperparameters are sampled during the validation; $\rho$, $\sigma_{in}$, $\sigma_{f}$, $\sigma_{\beta}$, $\sigma_{\tau}$, $\alpha$ and $\lambda$ are sampled uniformly on a $\log_{10}$ scale, while $k_{\beta}$ and $k_{\tau}$ are sampled uniformly.
@@ VI.2 Prediction performance
Figure 2: Short-term prediction performance of (a) standard parameter-aware ESN (9), (b) advection-augmented parameter-aware ESN (9), (c) thermoacoustic ESN (21). Relative $\ell_{2}$ error between prediction and ground truth of $[\mathbf{\bm{\eta}};\;\mathbf{\bm{\mu}}]$ over 20 time units on a $\log_{10}$ scale.
@@
The computed adjoint sensitivities are only as accurate as the model we have. Therefore, the first task of the ESN is to accurately emulate the dynamics of regimes both seen and unseen during training. This includes two subtasks: (i) time-accurate prediction from a given initial condition, and (ii) convergence to the correct attractor. The latter is necessary for long-term performance because we are interested in the sensitivity of long-time averaged quantities. Figure 2 shows the short-term prediction performance of the standard, advection-augmented and thermoacoustic ESNs over a grid of $\beta$ and $\tau$ values. For a given regime, after a washout stage, the ESN is run in closed-loop for prediction. This means that the network is forecasting the dynamics. The relative $\ell_{2}$ error (27) between prediction and ground truth of $[\mathbf{\bm{\eta}};\;\mathbf{\bm{\mu}}]$ over 20 time units is averaged over 5 folds starting from random time steps, and 5 realisations of the ESN. Standard parameter-aware ESN is not capable of sustaining the limit cycle solutions for approximately 10 periods even for the training regimes, whereas the augmentation of the advection variables $\mathbf{v}$ in the input and output significantly improves the performance. Notably, the T-ESN, generalizes to unseen regimes. Owing to the nonlinearity in the system, the prediction performance decreases as we move away from the training regimes.
@@
In Fig. 3, we show the performance of the T-ESN on (a) a limit cycle regime seen during training ($\beta=2.0,\;\tau=0.25$), and (b) an unseen limit cycle regime ($\beta=4.5,\;\tau=0.12$). Based on Fig. 2, we choose regime (b) as an example of an unseen regime with higher relative $\ell_{2}$ error compared to the other limit cycle regimes. In these figures, we assess the performance by (from left to right) short-term time-series prediction for 20 time units after washout, the probability density function (PDF), i.e., statistics, and amplitude spectrum of the first pressure mode $\mu_{1}$, and the acoustic energy $E_{ac}$ (20). For the PDF and amplitude spectrum, we feed the initial condition of zeros except $\eta_{1}(t_{0})=1.0$ in the washout stage, let the ESN evolve in closed-loop, discard a transient of 200 time units and use the next 1000 time units. This procedure and results verify that the ESN has learned the correct limit cycle attractor because it converges to this set of states from an initial condition that is not in the training set.
Refer to captionRefer to caption
@@
Figure 3: Thermoacoustic Echo State Network learns limit cycle regimes both seen and unseen during training. Short-term prediction, probability distribution function (PDF), and amplitude spectrum of the first pressure mode and acoustic energy of (a-f) a limit cycle regime seen during training ($\beta=2.0,\;\tau=0.25$) and (g-l) an unseen limit cycle regime ($\beta=4.5,\;\tau=0.12$). In the PDFs, the mean of an ensemble of 5 realisations of the ESN is shown, one standard deviation from the mean is plotted in lighter colour.

### v2

@@ III Adjoint of the Echo State Network
In this section, we derive the adjoint of the parameter-aware ESN. In closed-loop, the ESN is governed by,
$\mathbf{r}(i+1)=(1-\alpha)\mathbf{r}(i)+\alpha\tanh(\mathbf{W}_{in}[\mathbf{W}_{out}\mathbf{r}(i);\mathrm{diag}(\mathbf{\bm{\sigma}}_{p})(\mathbf{p}-\mathbf{k}_{p})]+\mathbf{W}\mathbf{r}(i)),$ (10)
which is an autonomous dynamical system. This is the configuration that we use for the computation of the adjoint. We first tackle sensitivity to the system’s parameters, and then to the system’s initial conditions.
III.1 Sensitivity to the system’s parameters
We consider the time-averaged objective functional (2), which we wish to optimize. In discrete time, we reformulate it as a sum over a discrete number of time steps, denoted by $N$
@@ III.1 Sensitivity to the system’s parameters
$\mathbf{q}^{+}(i)$ $=\frac{1}{N}\frac{{\color[rgb]{0,0,0}d}\tilde{\mathcal{J}}(\mathbf{r}(i))}{{\color[rgb]{0,0,0}d}\mathbf{r}(i)}^{\top}+\frac{\partial\mathbf{r}(i+1)}{\partial\mathbf{r}(i)}^{\top}\mathbf{q}^{+}(i+1),$ (13b)
$\mathbf{q}^{+}(N)$ $=\frac{1}{N}\frac{{\color[rgb]{0,0,0}d}\tilde{\mathcal{J}}(\mathbf{r}(N))}{{\color[rgb]{0,0,0}d}\mathbf{r}(N)}^{\top}.$ (13c)
Equation (13) is in a form similar to the continuous time equations (5), however, we have replaced the system’s Jacobian with the ESN’s Jacobian, thus, the adjoint variables are the adjoint of the ESN’s reservoir state. Algorithmically, we first obtain the direct solution, $\mathbf{r}(i)$, as a baseline trajectory by running the ESN in closed-loop for a time window. Then, the adjoint equations (13b) are solved backwards in time starting from the terminal condition $\mathbf{q}^{+}(N)$ (13c). This procedure requires the computation of the Jacobian, $\partial\mathbf{r}(i+1)/\mathbf{r}(i)$, and the partial derivative with respect to the parameters $\partial\mathbf{r}(i)/\partial\mathbf{p}$, associated with the ESN. These gradients are provided in Appendix B.
@@
$\frac{{\color[rgb]{0,0,0}d}\tilde{\mathcal{J}}(\mathbf{r}(i))}{{\color[rgb]{0,0,0}d}\mathbf{r}(i)}=2\mathbf{W}_{out}^{\top}\mathbf{W}_{out}\mathbf{r}(i).$ (15)
@@
III.2 Sensitivity to the system’s initial conditions
We consider the sensitivity of an objective functional of the final reservoir state after $N$ time steps, i.e., $\mathcal{J}=\mathcal{J}(\mathbf{r}(N))$ to the initial conditions. In the ESN’s equations (9), the system’s initial conditions $\mathbf{y}(0)$ appear in the input as $\mathbf{y}_{in}(0)$. When computing the sensitivity to the initial input vector $\mathbf{y}_{in}(0)$, we need to make a distinction between whether the first step of the propagation of the ESN’s equations is performed in open-loop or closed-loop. Because of the recurrence in the ESN, the full state of the ESN is described by the reservoir state. Therefore, for the ESN’s estimate of the sensitivity to match the original system, we first run a washout so that the output of the reservoir synchronizes with the original system and then perform the first step in closed-loop. (For a use case of the resulting open-loop Jacobian, see an application in data assimilation in [25].) The sensitivity to the initial condition is
$\frac{d\mathcal{J}(\mathbf{r}(N))}{d\mathbf{y}_{in}(0)}=\frac{d\mathcal{J}(\mathbf{r}(N))}{d\mathbf{r}(N)}\frac{d\mathbf{r}(N)}{d\mathbf{r}(1)}\left(\frac{\partial\mathbf{r}(1)}{\partial\mathbf{y}_{in}(0)}+\frac{\partial\mathbf{r}(1)}{\partial\mathbf{r}(0)}\frac{d\mathbf{r}(0)}{d\mathbf{y}_{in}(0)}\right),$ (16)
where the gradients $\partial\mathbf{r}(1)/\partial\mathbf{y}_{in}(0)$ and $\partial\mathbf{r}(1)/\partial\mathbf{r}(0)$ are found by differentiating the open-loop equation (9) and the gradients $\partial\mathbf{r}(i+1)/\partial\mathbf{r}(i),\;i=1,2,\dots,N-1$ are found by differentiating the closed-loop equation (10). From the output of the ESN (6b), we find the gradient $d\mathbf{r}(0)/d\mathbf{y}_{in}(0)=\mathbf{W}_{out}^{\dagger}$, where $(.)^{\dagger}$ indicates pseudo-inverse. Alternatively, we compute $d\mathcal{J}(\mathbf{r}(N))/d\mathbf{r}(0)$ in closed-loop configuration for all time steps and then determine by the chain rule
$\frac{d\mathcal{J}(\mathbf{r}(N))}{d\mathbf{y}_{in}(0)}=\frac{d\mathcal{J}(\mathbf{r}(N))}{d\mathbf{r}(0)}\mathbf{W}_{out}^{\dagger}.$ (17)
The tangent linear equations are given by
$\frac{d\mathcal{J}(\mathbf{r}(N))}{d\mathbf{r}(0)}$ $=\frac{d\mathcal{J}(\mathbf{r}(N))}{d\mathbf{r}(N)}\mathbf{Q}(N),$ (18a)
$\mathbf{Q}(i+1)$ $=\frac{\partial\mathbf{r}(i+1)}{\partial\mathbf{r}(i)}\mathbf{Q}(i),$ (18b)
$\mathbf{Q}(0)$ $=\mathbf{I},$ (18c)
where $\mathbf{Q}(i)\equiv d\mathbf{r}(i)/d\mathbf{r}(0),\;\mathbf{Q}(i)\in\mathbb{R}^{n_{r}\times n_{r}}$ and $\mathbf{I}$ is the identity matrix. The adjoint equations are
$\frac{d\mathcal{J}(\mathbf{r}(N))}{d\mathbf{r}(0)}$ $=\mathbf{q}^{+\top}(0),$ (19a)
$\mathbf{q}^{+}(i)$ $=\frac{\partial\mathbf{r}(i+1)}{\partial\mathbf{r}(i)}^{\top}\mathbf{q}^{+}(i+1),$ (19b)
$\mathbf{q}^{+}(N)$ $=\frac{d\mathcal{J}(\mathbf{r}(N))}{d\mathbf{r}(N)}^{\top}.$ (19c)
In this case, the number of “parameters” is equal to the dimension of the reservoir state vector, which is typically large, and therefore the adjoint approach significantly reduces the computational cost as compared to finite differences or the tangent linear approach.
@@ IV Thermoacoustics
We consider a prototypical thermoacoustic system known as the Rijke tube, which consists of a duct with a heat source located inside. The Rijke tube is a setup used to propose fundamental methods of thermoacoustic instabilities because it captures the qualitative nonlinear dynamics and bifurcations observed in real-life applications [e.g., 43, 33, 6]. In this model, we make the following assumptions; (i) the acoustics are one-dimensional, (ii) the mean-flow has a low Mach number and the acoustic variables are modelled as perturbations on top of a mean flow with uniform density across the tube, (iii) the heat release is compact, and (iv) the boundary conditions are ideal, i.e., fully reflective. The governing partial differential equations (PDEs) are derived from conservation of momentum, energy, and mass [e.g., 5]
@@ V Thermoacoustic Echo State Network
$\frac{d\mathcal{J}}{d\mathbf{p}}$ $={\color[rgb]{0,0,0}\frac{1}{N}}\sum_{i=1}^{N}\frac{{\color[rgb]{0,0,0}d}\tilde{\mathcal{J}}(\mathbf{r}(i))}{{\color[rgb]{0,0,0}d}\mathbf{r}(i)}\mathbf{Q}(i),$ (27a)
@@
$\mathbf{q}^{+}(i)$ $=\frac{1}{N}\frac{{\color[rgb]{0,0,0}d}\tilde{\mathcal{J}}(\mathbf{r}(i))}{{\color[rgb]{0,0,0}d}\mathbf{r}(i)}^{\top}+\frac{\partial\mathbf{r}(i+1)}{\partial\mathbf{r}(i)}^{\top}\mathbf{q}^{+}(i+1)+\frac{\partial\mathbf{r}(i+1+N_{\tau})}{\partial\mathbf{r}(i)}^{\top}\mathbf{q}^{+}(i+1+N_{\tau}),$ (28b)
$\mathbf{q}^{+}(N)$ $=\frac{1}{N}\frac{{\color[rgb]{0,0,0}d}\tilde{\mathcal{J}}(\mathbf{r}(N))}{{\color[rgb]{0,0,0}d}\mathbf{r}(N)}^{\top},$ (28c)
@@
The derivation of the time-delayed adjoint is provided in Appendix A. The time-delayed Jacobian is provided in Appendix B. (The tangent linear and adjoint equations for the initial condition sensitivity are found similarly by including the time-delayed or time-advanced Jacobian to the Eqs. (18) and (19) in Sec. III, respectively.)
@@
VI Learning thermoacoustics and parameter sensitivities
@@ VI.1 Training and validation
For training and validation, we create a dataset from time-delays, $\tau=\{0.1,0.15,0.2,0.25,0.3\}$ and heat release strengths, $\beta=\{1.0,2.0,3.0,4.0,5.0\}$, giving a total of 25 thermoacoustic regimes. These parameters generate limit cycle solutions except for $(\beta=5.0,\tau=0.15)$, which produces a quasiperiodic solution. We fix the flame location $x_{f}=0.2$, modal damping coefficients $c_{1}=0.1$ and $c_{2}=0.06$, the number of Galerkin modes $n_{g}=4$, and the number of Chebyshev discretization $n_{c}=10$ [6]. With four Galerkin modes, we observe the rich nonlinear behaviour associated with thermoacoustics. We generate the data by time-marching the ODEs (22) using the scipy.odeint tool with a time step of 0.001, which we downsample to 0.01 for the ESN. We remove a transient of 200 time units, select the next 4 time units as washout and 8 time units as training data. (We found that increasing the training length further does not improve the performance whereas it increases the computational effort.) We randomly split this dataset into 20 regimes for training and 5 regimes for validation of the hyperparameters. During validation, we evaluate the closed-loop performance of 5 random realisations of the ESN, i.e., realisations of the randomly generated input and state matrices, over 5 trajectories, or folds, of length 4 time units (approximately 2 periods) starting from random time steps. As error metric, we compute relative $\ell_{2}$-error between the prediction, $\hat{\mathbf{y}}$ and the ground truth $\mathbf{y}$
@@
We consider three networks: (i) the standard parameter-aware ESN (9) with the Galerkin amplitudes as the input and output vectors $\mathbf{y}_{in}(i)=\hat{\mathbf{y}}(i)=[\hat{\mathbf{\bm{\eta}}}(i);\;\hat{\mathbf{\bm{\mu}}}(i)]$, (ii) we augment the input and output vectors with the advection variable $\mathbf{y}_{in}(i)=\hat{\mathbf{y}}(i)=[\hat{\mathbf{\bm{\eta}}}(i);\;\hat{\mathbf{\bm{\mu}}}(i);\;\hat{\mathbf{v}}(i)]$, and (iii) thermoacoustic ESN (T-ESN) (26). The standard and the advection-augmented parameter-aware ESNs have both $\beta$ and $\tau$ as explicit parameters in the input channel, i.e., in Eq. (9) $\mathbf{p}=[\beta;\;\tau]$. The T-ESN has $\beta$ as a parameter in the input channel and $\tau$ is modelled as a time-delay in the delayed velocity term in the input channel $u_{f}(t-\tau)$ (26a). For comparison, we fix the reservoir size, i.e., number of reservoir state variables $n_{r}$, to 1200 and connectivity of the reservoir state variables, i.e., average number of nonzero elements per row in the state matrix $\mathbf{W}$, to 20. The optimal hyperparameters are provided in Table 1. The first row of Table 1 shows the ranges from which the hyperparameters are sampled during the validation; $\rho$, $\sigma_{in}$, $\sigma_{f}$, $\sigma_{\beta}$, $\sigma_{\tau}$, $\alpha$ and $\lambda$ are sampled uniformly on a $\log_{10}$ scale, while $k_{\beta}$ and $k_{\tau}$ are sampled uniformly.
@@ VI.2 Prediction performance
Figure 2: Short-term closed-loop prediction performance of (a) standard parameter-aware ESN (9), (b) advection-augmented parameter-aware ESN (9), (c) thermoacoustic ESN (26). Relative $\ell_{2}$ error between prediction and ground truth of $[\mathbf{\bm{\eta}};\;\mathbf{\bm{\mu}}]$ over 20 time units on a $\log_{10}$ scale.
@@
The computed adjoint sensitivities are only as accurate as the model we have. Therefore, the first task of the ESN is to accurately emulate the dynamics of regimes both seen and unseen during training. This includes two subtasks: (i) time-accurate prediction from a given initial condition, and (ii) convergence to the correct attractor. The latter is necessary for long-term performance because we are interested in the sensitivity of long-time averaged quantities. Figure 2 shows the short-term prediction performance of the standard, advection-augmented and thermoacoustic ESNs over a grid of $\beta$ and $\tau$ values. For a given regime, after a washout stage, the ESN is run in closed-loop for prediction. This means that the network is forecasting the dynamics. The relative $\ell_{2}$ error (32) between prediction and ground truth of $[\mathbf{\bm{\eta}};\;\mathbf{\bm{\mu}}]$ over 20 time units is averaged over 5 folds starting from random time steps, and 5 realisations of the ESN. Standard parameter-aware ESN is not capable of sustaining the limit cycle solutions for approximately 10 periods even for the training regimes, whereas the augmentation of the advection variables $\mathbf{v}$ in the input and output significantly improves the performance. The T-ESN generalizes to unseen regimes. Owing to the nonlinearity in the system, the prediction performance decreases as we move away from the training regimes.
@@
In Fig. 3, we show the performance of the T-ESN on (i) a limit cycle regime seen during training ($\beta=2.0,\;\tau=0.25$), and (ii) an unseen limit cycle regime ($\beta=4.5,\;\tau=0.12$). Based on Fig. 2, we choose regime (ii) as an example of an unseen regime with higher relative $\ell_{2}$ error compared to the other limit cycle regimes. In these figures, we assess the performance by (from left to right) short-term time-series prediction for 20 time units after washout, the probability density function (PDF), i.e., statistics, and amplitude spectrum of the acoustic energy $E_{ac}$ (25). For the PDF and amplitude spectrum, we feed the initial condition of zeros except $\eta_{1}(t_{0})=1.0$ in the washout stage, let the ESN evolve in closed-loop, discard a transient of 200 time units and use the next 1000 time units. This procedure and results verify that the ESN has learned the correct limit cycle attractor because it converges to this set of states from an initial condition that is not in the training set.
@@
Figure 3: Thermoacoustic Echo State Network learns limit cycle regimes both seen and unseen during training. Short-term prediction, probability distribution function (PDF), and amplitude spectrum of the acoustic energy of (a-c) a limit cycle regime seen during training ($\beta=2.0,\;\tau=0.25$) and (d-f) an unseen limit cycle regime ($\beta=4.5,\;\tau=0.12$). In the PDFs, the mean of an ensemble of 5 realisations of the ESN is shown, one standard deviation from the mean is plotted in lighter colour.